What are the empirical and molecular formulas of a compound that is 40.27% potassium, 26.78% chromium, and 32.95% oxygen? The molar mass of the compound is 194.2 g/mol.

1
Expert's answer
2022-02-14T23:36:03-0500

Calculating the masses;

(40.27/100)× 194.2= K= 78.2g

(26.78/100)× 194.2= Cr= 52g

(32.95/100)× 194.2= O= 64g

Calculating the mole ratio_K:Cr:O

Moles= mass/rfm

78.2/39= K= 2

52/52= Cr= 1

64/16= O= 4

Formula= K2CrO4